# CataractBot: An LLM-Powered Expert-in-the-Loop Chatbot for Cataract Patients

### v1

@@ 4.2. Procedure for Doctors and Coordinators
In addition to recruiting patients and attendants, we recruited four doctors, each regularly performing cataract surgeries. Among them, three served as operating doctors, one as an escalation doctor and KB Update expert. Additionally, two patient coordinators were recruited, both specializing in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator. The escalation doctor, escalation coordinator, and KB update expert held senior positions in the hospital and carried administrative responsibilities. The KB update expert is part of the research team, and was actively involved in curating the knowledge base. This expert played a pivotal role in recruiting all other doctors and coordinators individually through one-on-one discussions.
To ensure preparedness, all participating doctors and coordinators attended one of two one-hour training sessions conducted by three researchers. During these sessions, CataractBot was demonstrated, and participants were onboarded onto the platform. The researchers assumed patient roles, asking medical and logistical questions for the doctors and coordinators to verify and edit. The first session revealed several bugs, and feedback from doctors and coordinators was collected to enhance the bot. For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in tagged WhatsApp messages (Figure 1E).
After using CataractBot for two weeks or longer, the participating doctors and coordinators were interviewed by the first author. These interviews followed a semi-structured format, focusing on the bot’s usability, its integration into their workflow, and its impact on interactions with patients. The interviews were audio-recorded, conducted in English, in-person, and typically lasted 45-60 minutes. None of the participants—patients, attendants, doctors, or coordinators—received any financial incentives for their participation.
@@ 4.3. Participants
A total of 31 patient-attendant pairs (Table 1), consisting of 19 patients and 30 attendants, took part in our study. Due to the predominantly elderly demographic undergoing cataract surgery, 12 patients lacked access to WhatsApp/smartphones and were unable to participate. One patient was not accompanied by an attendant. The average age of the 19 participating patients (3 females) was 58.88.01 years. They exhibited a diverse distribution in preferred languages (5 English, 5 Hindi, 5 Kannada, 3 Tamil, and 1 Telugu) and highest education levels (6 Grade 10, 4 Grade 12, 6 Bachelors, and 3 Masters). In contrast, the attendants were younger, predominantly fluent in English, and well-educated. The average age of the 30 participating attendants (11 female) was 37.210.3 years. A majority (23) preferred English, with 2 each favoring Hindi, Kannada, and Tamil, and 1 preferring Telugu. Education levels were mostly high, with 10 having Masters and 13 having Bachelors degrees. Among the 31 patient-attendant pairs, 22 participated in the surgery-day interview, and 10 took part in the one-week post-surgery interview. Various factors influenced participation, including time constraints (resulting in a few declined interviews) and data reaching saturation. Additionally, four doctors (three as operating doctors, one as escalation doctor, and one as KB update expert) and two patient coordinators (one as coordinator and one as escalation coordinator) participated in the study (Table 2). Note: The roles of escalation doctor and KB update expert were performed by the same doctor.
@@ 4.4. Data collection and analysis
The first author translated and transcribed all interviews (totalling 11.5 hours) with patients, attendants, doctors and coordinators, into English soon after they were conducted. Throughout the study, all authors regularly engaged in discussions to review observations and recent interviews, fostering a nuanced understanding of bot usage and practices while identifying emerging themes of interest. For the analysis of interview transcripts, we employed an inductive coding approach, with the first author open-coding (Williams and Moser, 2019) all interviews line-by-line. Subsequently, two authors collaboratively reviewed these codes to identify interesting themes in the data (similar to (Okolo et al., 2021)). We iterated on these themes to distill higher-level themes (like “Usage”, “Features”, and “Privacy and Trust”) that we present in our findings below.
Note: While there were a total of 49 patients and attendants using CataractBot, at times, we treat them as 31 participants (patient-attendant pairs). This choice stems from instances highlighted during interviews, where patients instructed attendants to query the bot using their phones. Additionally, both were interviewed together. Only when comparing their distinct roles, we analyzed their usage log data separately.
@@ 4.5. Positionality
All authors are of Indian origin, and currently reside in Bengaluru, India. Among them, two identify as female and the rest as male. One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly. Their involvement spanned designing the study, shaping research questions, providing feedback on initial CataractBot versions, contributing to knowledge base curation, and serving as the KB update expert. Another author is affiliated with the Quality Control and Patient Safety team, who ensured that any risk of harm to the patients was mitigated. Two authors specialize in HCI and design, three possess healthcare research experience, and four have expertise in software development. We approached this research drawing on our individual experiences and learning from working at the intersection of computing and healthcare in the Indian context. Our interest in making healthcare more accessible to the masses has informed the design of CataractBot, guided our study design, and shaped our data analysis.
@@ 5. Findings
Overall, 31 patient-attendant222We use the terms ‘participants’ or ‘information-seekers’ to refer to patients and their attendants, and ‘experts’ to refer to doctors and patient coordinators. pairs sent 343 messages to CataractBot, with an average of 11.068.44 messages per pair. Among the 19 participating patients, only 12 interacted with the bot by sending one or more messages, while all 30 participating attendants, except one, messaged the bot. Patients sent a total of 89 messages (7.45.0 message/patient), while attendants sent 254 messages (8.86.8 message/attendant). Out of the 343 messages, 225 were medical queries, 87 logistical questions, 27 small talk, and 4 others. Doctors directly approved 69.8% (157) of the bot-generated answers to medical questions and provided corrections for 29.3% (66) of the answers, while patient coordinators directly approved 58.6% (51) of logistical answers and provided corrections for 40.2% (35) of the answers. Note: Only two medical questions and one logistical question did not receive any response from the experts, while for eight questions (five medical and three logistical) an expert responded with ‘No’ to the poll without providing a correction. Below, in our findings, we delve into how, why, and when participants used CataractBot, examine the impact of novel features, and conclude by addressing privacy and trust concerns in the context of a chatbot in a medical use case.
@@ 5.1. CataractBot Usage
5.1.1. WHY
Patients and attendants used CataractBot for various reasons, such as seeking clarification on answers provided by the expert, addressing questions they had forgotten to ask the doctor or coordinator, inquiring about matters they felt less comfortable asking the expert directly, resolving emergency situations, and obtaining status updates. In the usability survey, 77.3% of participants found CataractBot’s responses to be useful, appropriate, and informative, while 91.0% said they would be willing to use it (or a similar bot) in the future.
Ten participants highlighted the bot’s helpfulness in recalling information provided by the doctor/coordinator during in-person consultations, either due to forgetfulness or incomplete understanding. As A13 stated: “We are human minds… cannot give complete attention all the time.” Patients found comfort in “written” information, ensuring persistence and enabling them to refer back to it when needed, “I need not remember everything.” (A20). The value of persistent information in the context of a text-based bot has been underscored in prior research (Jain et al., 2018). Despite knowing the answers to certain questions, six participants mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17).
Doctors highlighted that many patients/attendants now come prepared for face-to-face conversations, optimizing their limited time with experts and reducing multiple visits to the hospital: “When you’re going to a surgery… there are ten things to ask. And when you are in front (of the doctor), you’ll remember only two things. Most of the patients nowadays come with a list of questions.” (D1). CataractBot has the potential to alleviate this burden of remembering every important question. Also, this flexibility of using CataractBot at one’s convenience–“as and when I get anything on my mind to ask” (P17)–was noted. The asynchronous nature of the conversation is beneficial, as information seekers do not have to constantly wait, either physically or over a call, for an expert’s response. A majority of participants (86.4%) agreed that the bot responded quickly. On average, the bot took 11.810.2 seconds to respond. Expert verification took 158.7168.4 minutes; we found verification without edits (151.1171.4 minutes) was faster than verification with edits (174.5161.4 minutes) which followed by 191.3172.4 minutes to provide the correction for the bot. This asynchronous feature not only benefited patients but also motivated experts to engage with the bot. The minimal time required to verify and edit an answer is another motivating factor.
“I can do it anytime… I do not have to think about phrasing and grammar… The bot is going to make it simpler and readable.” – D2.
Three participants noted the challenge of obtaining “precise information” from external sources like Google, prompting them to use the bot. Extracting accurate and relevant information from Google search results and formulating effective search queries is non-trivial (ter Hofstede et al., 1996), particularly for individuals with low (tech-)literacy. Moreover, due to existing power differences between doctors and patients (Joseph-Williams et al., 2014), such individuals often hesitate to ask questions. Experts acknowledged that educated individuals have access to alternative information sources, and are comfortable posing questions. In contrast, “less educated, local language speaking patients” are the ones for whom CataractBot “can do wonders” (D3).
“They have lots of doubts… even if I talk to them in their native language, they hesitate to ask… because it sounds silly… Could be because of the lower socio-economic background, or education.” – D1.
@@ 5.1.1. WHY
Five participants, including three with Grade-12 education, exclusively relied on CataractBot for information (asking 57 questions in total) as no one in their social circles had experience with cataract surgery, and they were not accustomed to seeking information online. Another expressed that inquiring with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgemental nature:
“(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” – A4.
Although explicitly mentioned during onboarding that CataractBot should not be used during emergencies, three patients/attendants attempted to use it to address emergencies, such as allergic reactions and taking the wrong medications. This raised concerns among two experts as CataractBot is not equipped to diagnose or handle emergencies. Expert responses are not immediate, and patient queries may lack the necessary details for the expert to provide an appropriate response without further clarification.
“If the patient has a little redness… and is supposed to take a flight… the bot may say “use the same medication no need to worry”, or it says “it’s an emergency come to the hospital, cancel the flight… In either scenario, it’s going to be difficult, because you don’t know what the redness is due to.” – D1.
The doctor suggested that the bot should be capable of detecting emergency situations, refrain from providing a response, and instead advise users to call the hospital.
Seven participants emphasized that the bot saved them time, as it eliminated the need to visit or call the hospital. Traveling to the hospital can be time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as A4 stated:
“Yesterday she (coordinator) called… she told to come tomorrow, that’s all. I was about to ask a few questions but then she cut the call. Then I called back. It was busy, may be because she was talking to others.” – A4.
@@
Despite the compelling reasons for using the bot, a few participants used it minimally. Out of the 49 participants, 9 did not use the bot at all, while 7 sent 1-2 messages to the bot. Six participants mentioned that their in-person interactions with the doctor and patient coordinator at Sankara Eye Hospital sufficed to address their queries, reducing their reliance on the bot. Two participants considered cataract surgery to be a “simple” procedure, leading them to have no questions. This perception could stem from their previous experience with cataract surgery, undertaken six months or more ago for their other eye. The patient coordinator added that abundant cataract surgery related information is readily available online, unlike specialised surgeries. This suggests that CataractBot has the potential to play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing patients’ information needs or where power imbalances make individuals less likely to seek answers. Additionally, a similar bot could prove valuable for more complex (surgical) procedures.
@@
Participants also complained about receiving “I don’t know” (IDK) responses (57 total) and/or experiencing delays in responses. A total of 36 (10.5%) messages pertained to requests for status updates—seeking information about the surgery time and discharge time. Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, lacking access to patient details and resulting in IDK responses for the non-generic questions. In such instances, the patient coordinator later provided appropriate responses. While participants acknowledged the appropriateness of schedule-related IDK responses due to the lack of access to patient details, for other queries, they suggested providing a “tentative response” rather than a blank IDK. For example, for the question “Mine is general anaesthesia or local anaesthesia?” (A10), the bot provided an IDK response, adhering to the cautious approach advised by hospital experts due to medical and legal concerns. However, A10 wanted a response such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”. From an expert’s perspective, responding to certain IDK questions was challenging as they lacked the full context. E.g., A21 asked a complex, personalized, and high-stakes question: “I want to postpone the surgery to February… is this okay or do I need to get this operated immediately?” In response, the doctor corrected the bot’s IDK response with: “You come and check.”. This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye care check; the doctor would know her condition… I don’t want to come and visit the doctor [again]… it will waste my 2-3 hours.” (A21).
@@
For time-sensitive questions, A1 and A9 desired “quick response” from the bot. The delay in expert verification led them to perceive the bot as ineffective. “The dosage… every 2 hours take 1 drop. If it takes anything more than 2 hours to respond [to my eye drop related question], the purpose is already defeated… If it is not providing me timely help… as a consumer, I don’t see value add.” (A1). This was reflected in the usability survey as well, with 27.2% of participants selecting ‘Disagree’ or ‘Strongly Disagree’ regarding the bot’s grasp of medical knowledge.
Experts used CataractBot driven by their sincere commitment to help patients. As D2 put it “I want to help my patients… I am least bothered about the timing. If he is having a problem, he should be answered… after all, we are living for the patient’s sake.”. Moreover, the asynchronous nature of the bot provided experts with the flexibility to offer verifications at their convenience. Finally, experts appreciated the intermediary role of the bot. The bot-in-the-middle model introduced a layer of separation from patients: “From a doctor’s standpoint, it is very good because it puts in a curtain between us.” (D4).
5.1.2. HOW
The workflow employed by our patients and attendants varied, with time-sensitivity being a crucial factor. For time-insensitive questions, participants tended to consume the unverified response, wait for the expert’s verification, and later, consume the (updated) verified response. Our participants demonstrated a clear understanding of the bot’s workflow and expressed being “okay to wait” since the response was being verified by “busy doctors”.
@@ 5.1.2. HOW
“I know that it will take its own time… as its doctors… based on their busy schedule, their availability… It’s not like there’s some dedicated doctor monitoring it all the time, right? If that is the case, then bot is not required. They can directly reply.” – A17.
@@
A few participants mentioned that they hardly noticed the wait as they were occupied with their day-to-day activities. They would ask a question, leave their phones, and check them later when free. In such cases, notifications played a key role in the information-seeking workflow. The notification message (Figure 1G), stating that the expert had provided a verified response, made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory. For questions where participants received an unsatisfactory IDK response, they sought information from other sources, including Google and asking friends and family.

### v2

@@ 5.2. Procedure for Doctors and Coordinators
In addition to patients and attendants, we recruited four doctors and two patient coordinators. Among the doctors, three served as operating doctors, and one as both an escalation doctor and KB update expert. Both patient coordinators specialize in handling cataract surgery patients, with one serving as a coordinator and the other as an escalation coordinator. The escalation doctor and escalation coordinator held senior positions in the hospital and carried administrative responsibilities. To ensure preparedness, all participating experts attended one of two one-hour training sessions conducted by three researchers. During these sessions, CataractBot was demonstrated, and they were onboarded onto the platform. The researchers assumed patient roles, asking medical and logistical questions for the doctors and coordinators to verify and edit. The first session revealed a few bugs, and feedback was collected to enhance the bot. For example, a doctor suggested displaying the question first (rather than the patient’s demographics) in verification messages, as only the top portion is visible in ‘tagged’ WhatsApp messages (Figure 1E). After using CataractBot for two weeks or longer, the participating experts were interviewed by the first author. These interviews followed a semi-structured format, focusing on the bot’s usability, its integration into their workflow, and its impact on interactions with patients. The interviews were audio-recorded, conducted in English, in-person, and typically lasted 45-60 minutes.
@@ 5.3. Participants
A total of 31 patient-attendant pairs (Table 1), consisting of 19 patients and 30 attendants, took part in our study. Due to the predominantly elderly demographic undergoing cataract surgery, 12 patients lacked access to WhatsApp/smartphones and were unable to participate but their attendants used the bot. One patient was not accompanied by an attendant. The average age of the 19 participating patients (3 females) was 58.88.01 years. Although females comprised 32.7% of our study participants, only 15.8% of the patients were female. This aligns with prior work showing a gender disparity in accessing cataract surgery in India (Prasad et al., 2020; Ye et al., 2020). Patients had diverse preferred languages (5 English, 5 Hindi, 5 Kannada, 3 Tamil, and 1 Telugu) and education levels (6 Grade 10, 4 Grade 12, 6 Bachelors, and 3 Masters). In contrast, the attendants were younger, predominantly fluent in English, and well-educated. The average age of the 30 participating attendants (11 female) was 37.210.3 years. A majority (23) preferred English, with 2 each favoring Hindi, Kannada, and Tamil, and 1 preferring Telugu. Education levels were mostly high, with 10 having Masters and 13 having Bachelors degrees. Among the 31 patient-attendant pairs, 22 participated in the surgery-day interview, and 10 took part in the one-week post-surgery interview. Various factors influenced participation, including time constraints (resulting in a few declined interviews) and data reaching saturation. Additionally, four doctors and two patient coordinators participated in the study (Table 2). Note: There was a partial overlap between participants in this study and the formative study, with two doctors and one coordinator participating in both.
@@ 5.4. Data collection and analysis
The first author translated and transcribed all interviews (totalling 11.5 hours) into English soon after they were conducted. Throughout the study, all authors regularly engaged in discussions to review observations and interviews. To analyze interview transcripts, we employed an inductive coding approach, with the first author open-coding (Williams and Moser, 2019) all interviews line-by-line. Subsequently, two authors collaboratively reviewed these codes to identify interesting themes in the data (similar to (Okolo et al., 2021)). We iterated on these themes to distill higher-level themes (like “Usage”, “Features”, and “Privacy and Trust”) that we present in our findings below. Note: While there were a total of 49 patients and attendants using CataractBot, at times, we treat them as 31 participants (patient-attendant pairs). This choice stems from instances highlighted during interviews, where patients instructed attendants to query the bot using their phones. Additionally, both were interviewed together. Only when comparing their distinct roles, we analyzed their usage log data separately.
@@ 5.5. Positionality
All authors are of Indian origin, and currently reside in Bangalore, India. Among them, two identify as female and the rest as male. One author is a practising ophthalmologist, performing over 10 cataract surgeries weekly, and another author is part of the Quality Control and Patient Safety team at the same hospital. Both were involved in designing the study, shaping research questions, providing feedback on initial bot versions, curating the KB, and ensuring that any risk of harm to the patients was mitigated. The ophthalmologist also served as the KB update expert. Two authors specialize in HCI and design, three possess healthcare research experience, and four have expertise in software development. We approached this research drawing on our individual experiences and learning from working at the intersection of computing and healthcare in the Indian context. Our interest in making healthcare more accessible to the masses has informed the design of CataractBot, guided our study design, and shaped our data analysis.
@@ 6. Findings
Overall, 31 patient-attendant pairs sent 343 messages to CataractBot, with an average of 11.068.44 messages per pair. Out of the 343 messages, 225 were medical questions, 87 logistical, 27 small talk, and 4 others. Doctors directly approved (said ‘Yes’ to) 69.8% (157) of the bot-generated answers to medical questions, while patient coordinators directly approved 58.6% (51) of logistical answers.In the usability survey, 77.3% of participants found CataractBot’s responses to be useful, appropriate, and informative, while 91.0% said they would be willing to use it in the future. Below, in our findings, we delve into how participants used CataractBot (answering RQ1 and RQ3), examine the impact of novel features introduced in CataractBot (answering RQ2), and explore privacy and trust implications of using an LLM-powered chatbot (answering RQ1).
@@ 6.1. CataractBot Usage
6.1.1. Reasons for (Not) Using CataractBot
Information seekers used CataractBot to address questions they had forgotten or were uncomfortable to ask the doctor or patient coordinator, clarifying answers, and seeking updates. Doctors stated that most information seekers now-a-days come prepared for face-to-face conversations, optimizing their limited time with experts and reducing multiple visits. CataractBot can alleviate the burden of remembering every question. Ten participants highlighted that the bot helped in recalling information provided by the doctor/coordinator during in-person consultations, which was either forgotten or incompletely understood. As A13 stated: “We are humans… cannot give complete attention all the time.” Patients found comfort in “written” information, ensuring persistence and enabling them to refer back to it when needed (similar to (Jain et al., 2018)), “I need not remember everything.” (A20).
@@ 6.1.1. Reasons for (Not) Using CataractBot
Due to existing power differences between doctors and patients (Joseph-Williams et al., 2014), individuals often hesitate to ask questions. Experts acknowledged that educated individuals have access to other information sources, and are comfortable posing questions. In contrast, “less educated, local language speaking patients” are the ones for whom CataractBot “can do wonders” (D3). Five participants, including three with Grade-12 education, exclusively relied on CataractBot for information (asking 57 questions in total) as no one in their social circles had experience with cataract surgery, and they were not accustomed to seeking information online. Another expressed that checking with the bot was more comfortable than approaching a doctor, citing the bot’s non-judgemental nature: “(The patient) has cholesterol issues and heart surgery… the doctor would make faces when I ask about food intake… but not the bot.” (A4). Six participants also mentioned using the bot to “double-check” (A7) and seek “reassurance” (A17), in spite of knowing certain answers.
The asynchronous feature was helpful, as the information seekers did not have to constantly wait, either physically or over a call, for an expert’s response. Seven participants emphasized that the bot saved them time, as it eliminated the need to visit or call the hospital. Traveling to the hospital can be time-consuming, and contacting busy hospitals over the phone is challenging due to high patient loads, as “the phone is always busy” (A4). Also, information seekers found the bot easy to use, with 86.2% rating ‘Agree’ or ‘Strongly Agree’ in response to the usability question (Appendix A.2), mainly due to their familiarity with WhatsApp. Additionally, features such as support for multiple languages, audio support, and related questions (Figure 2), lowered barriers to initiating conversation.
Experts used CataractBot driven by their commitment to help patients.
“I want to help my patients… I am least bothered about the timing. If he is having a problem, he should be answered… after all, we are living for the patient’s sake.” – D2.
Experts also appreciated the intermediary role of the bot, which introduced a layer of separation from patients: “From a doctor’s standpoint, it is very good because it puts in a curtain between us.” (D4).
@@
Despite the compelling reasons to use the bot, out of the 49 information seekers, 8 did not use the bot sending two or less messages. Six participants mentioned that their in-person interactions with the doctor and patient coordinator sufficed to address their queries, reducing their reliance on the bot. Two participants considered cataract surgery to be a “simple” procedure, leading them to have no questions. This perception stemmed from their previous experience with cataract surgery of their other eye. This suggests that CataractBot could play a more pivotal role in primary healthcare centers, where the human infrastructure may be less capable of addressing patients’ information needs or where power imbalances make individuals less likely to seek answers.
@@
Participants also complained about receiving “I don’t know” (IDK) responses (57 total) and experiencing delays in experts verifying responses. A total of 36 (10.5%) messages requested status updates—seeking information about the surgery time and discharge time. Due to privacy concerns, CataractBot was not integrated with the hospital’s patient management system, resulting in IDK responses for such questions. The patient coordinator later provided appropriate responses. While participants acknowledged the appropriateness of IDK responses due to the lack of access to patient details, they suggested providing a “tentative response” if possible. E.g., for the question “Mine is general anaesthesia or local anaesthesia?” (A10), the bot provided an IDK response. However, A10 expected a response such as “99% of cataract surgeries at Sankara Eye Hospital are done under local anesthesia”. From an expert’s perspective, responding to certain IDK questions was challenging as they lacked the full context. E.g., A21 asked: “I want to postpone the surgery to February… is this okay or do I need to get this operated immediately?” In response, the doctor corrected the bot’s IDK response with: “You come and check”. This correction frustrated the attendant, who stated: “But (the patient) already went through the complete eye care check; the doctor know her condition… I don’t want to come and visit the doctor [again]… it will waste 2-3 hours.” (A21). For time-sensitive questions, A1 and A9 desired “quick response” from the bot. The delay in expert verification led them to perceive the bot as ineffective.
@@
“The dosage… every 2 hours take 1 drop. If it takes anything more than 2 hours to respond [to my eye drop related question], the purpose is already defeated… If it is not providing me timely help… as a consumer, I don’t see value add.” – A1.
6.1.2. Workflow
Out of the total 343 messages sent by patients and attendants, all, except four, were sent between 7 am and 11 pm, with an average of 14.313.7 messages/hour. The messages were distributed throughout the day, as A18 mentioned: “Whenever I am idle at work or at home… whatever (question) comes to my mind, I just ask.” The workflow employed by participants varied, with time-sensitivity being a crucial factor. A majority of participants (86.4%) agreed that the bot responded quickly. On average, the bot took 11.810.2 seconds to respond. Expert verification took 162.9172.3 minutes. We found verification without edits (151.1147.1 minutes) to be significantly faster than verification with edits (191.3172.4 minutes) with t(265)=2.7, p¡0.05. For time-insensitive questions, participants were “okay to wait” for the expert’s verification, since the response was being verified by “busy doctors”. This demonstrated their clear understanding of the bot’s workflow.
@@ 6.1.2. Workflow
“I know that it will take its own time… as it’s doctors… based on their busy schedule, their availability… It’s not like there’s some dedicated doctor monitoring it all the time, right? If that is the case, then bot is not required. They can directly reply.” – A17.
@@
A few participants mentioned that they hardly noticed the wait as they were occupied with their day-to-day activities. They would ask a question, leave their phones, and check the answer later when free. In such cases, notifications played a key role as the notification message (Figure 1G), stating that the expert had provided a verified response, made the otherwise unnoticeable green tick (Figure 1F) prominent and the bot’s process self-explanatory.